# Revealing Language Model Trajectories via Kullback-Leibler Divergence

### v1

@@ 47 VAGOsolutions/Llama-3-SauerkrautLM-8b-Instruct llama-3 32 55400
48 lmsys/vicuna-13b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 40 48653
49 Qwen/Qwen2-1.5B Yang et al. (2024) others 28 46510
50 openlm-research/open_llama_7b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 32 44968
@@ 51 LinkSoul/Chinese-Llama-2-7b llama-2 32 40799
53 NousResearch/Hermes-2-Pro-Llama-3-8B "Teknium" et al. (2024b) llama-3 32 26797
54 lightblue/suzume-llama-3-8B-multilingual Devine (2024) llama-3 32 16442
55 abacusai/Llama-3-Smaug-8B Pal et al. (2024) llama-3 32 12873
@@ 56 NousResearch/Nous-Hermes-llama-2-7b llama-2 32 12019
58 elyza/ELYZA-japanese-Llama-2-7b-instruct Touvron et al. (2023b); Sasaki et al. (2023) llama-2 32 6296
59 togethercomputer/Llama-2-7B-32K-Instruct Liu et al. (2023) llama-2 32 5596

### v2

@@ 47 VAGOsolutions/Llama-3-SauerkrautLM-8b-Instruct llama-3 32 55400
48 lmsys/vicuna-13b-v1.5 arxiv:2307.09288; arxiv:2306.05685 llama-2 40 48653
49 Qwen/Qwen2-1.5B qwen2 others 28 46510
50 openlm-research/open_llama_7b together2023redpajama; arxiv:2302.13971; openlm2023openllama llama-1 32 44968
@@ 51 LinkSoul/Chinese-Llama-2-7b llama-2 32 40799
53 NousResearch/Hermes-2-Pro-Llama-3-8B Hermes-2-Pro-Llama-3-8B llama-3 32 26797
54 lightblue/suzume-llama-3-8B-multilingual arxiv:2405.12612 llama-3 32 16442
55 abacusai/Llama-3-Smaug-8B arxiv:2402.13228 llama-3 32 12873
@@ 56 NousResearch/Nous-Hermes-llama-2-7b llama-2 32 12019
58 elyza/ELYZA-japanese-Llama-2-7b-instruct arxiv:2307.09288; elyzallama2023 llama-2 32 6296
59 togethercomputer/Llama-2-7B-32K-Instruct arxiv:2307.03172 llama-2 32 5596